Your answer is partially correct. Try again Pharoah Company sells 9% bonds having a maturity value of $2,500,000 for $2,229,651. The bonds are dated January 1, 2017, and mature January 1, 2022 Interest is payable annually on January 1. Set up a schedule of interest expense and discount amortization under the straight-line method. (Round answers to 0 decimal places, e.g. 38,548.) Schedule of Discount Amortization Straight-Line Method Cash Paid Interest Expense Discount Amortized Carrying Amount of Bonds Year Jan. 1, 2017 Jan. 1, 2018 Jan. 1, 2019 Jan. 1, 2020 Jan. 1, 2021 Jan. 1, 2022 2229651 2229651 225000 54070 225000 54070 225000 54070 225000 54070 225000 54070 LINK TO TEXT
Answer #1

Straight line method :

Year Cash paid Interest expense Discount amortization Carrying amount of bonds
Jan 1,2017 2229651
Jan 1,2018 2500000*9% = 225000 (225000+54070) = 279070 (2500000-2229651/5) = 54070 2283721
Jan 1,2019 225000 279070 54070 2337791
Jan 1,2020 225000 279070 54070 2391861
Jan 1,2021 225000 279070 54070 2445931
Jan 1,2020 225000 279070 54070 2500001